lensth A soo nm i incident On the r/diamond interface shawn below. Calcul ete the e of ite recleeted and refracted mys ang 1.00 D, a mand, nzz 2.Ht b) e 4) For the Isht in pro blem 3, ate the following properti n weve ength in d D i z) speed of lisht in diamond and so the enerst of the ry in the dism and in electron volts b) 250 mm 207 2.Sev d) A 203 nm
Answer #1

(3)

Using Snell's law we have,

sinθi/sinθR = n2/n1 , where θi is the angle of incidence

or, sin450/sinθR = 2.42

or, θR = 170

Since, angle of incidence is equal to angle of refraction so,

θr = 450

correct option would be b)

***************************************************************************************************

4)

Speed of the light ray in air is,

v = 3X108 m/s

so frequency of the light is f = v/λ = (3X108 m/s)/(500X10-9m) = 6X1014 Hz

So speed in diamond is,

vd = v/n2  = (3X108 m/s)/(2.42)

or, vd = 1.24X108m/s

So wavelength in diamond is,

λd = vd/f = (1.24X108m/s)/(6X1014 Hz)

or, λd = 207 nm

So, the correct option is a).

P.S. - Amplitude of wave is necessary to calculate energy of the wave.

***************************************************************************************************
This concludes the answers. Check the answer and let me know if it's correct. If you need any more clarification, modification or correction, feel free to ask.....